(a) An ideal gas occupies a volume of 1.8 cm3 at 20°C and atmospheric pressure. Determine the number of molecules of gas in the container.

_____________ molecules

(b) If the pressure of the 1.8-cm3 volume is reduced to 2.4 ✕ 10−11 Pa (an extremely good vacuum) while the temperature remains constant, how many moles of gas remain in the container?
____________ mol
Answer #1

(a)

By ideal gas equation,

P*V = n*R*T

given,

Pressure = P = 1.013*10^5 Pa

volume(V) = 1.8 cm^3 = 1.8*10^-6 m^3

R = 8.314 J

Temperature(T) = 20 deg C = 293.15 K

n = moles of gas in container =??

So,

n = (P*V)/(R*T)

n = (1.013*10^5)*(1.8*10^-6)/(8.314*293.15)

n = 7.48*10^-5 mole

Since there are Na(6.022*10^23) molecules in 1 mole

So,

molecules of gas in n moles = n*Na = (7.48*10^-5)*(6.022*10^23)

molecules of gas = 4.5*10^19

(b.)

For value of pressure(P) = 2.4*10^-11 Pa

So,

now moles(n) = (2.4*10^-11)*(1.8*10^-6)/(8.314*293.15)

n = 1.77*10^-20

Therefore 1.77*10^-20 moles of gas remains in container.

Please upvote.